# Limits of open ASEP stationary measures near a boundary

### v1

@@ 2.1. Backgrounds
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $c=q^{j}As/\sqrt{t}$ and $d=1/\left(q^{j}A\sqrt{t}\right)<1$. We have $c/a=q^{j}s/t\notin\{q^{l}:l\in\mathbb{Z}\}$ since $s/t\in(q,1)$. Hence (2) holds.
@@
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $c=q^{j}C/\sqrt{t}$ and $d=s/(q^{j}C\sqrt{t})<1$. If $A<1$ then $a=A\sqrt{t}<1$ and (2) holds. If $C<1$ then $c=q^{j}C/\sqrt{t}<1$ and (2) holds. If $A,C\geq 1$ then $c/a=q^{j}C/(At)\notin\{q^{l}:l\in\mathbb{Z}\}$ by our assumption, hence (2) holds.
@@
Assume $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Choose the interval $I=[1,1+\varepsilon)$ from Theorem 2.5. Then for any $s\leq t$ in $I$ and any $x\in U_{s}$, the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is supported on $U_{t}$. In the high density phase, we have $P_{s,t}\left(y_{0}(s),\mathsf{d}y\right)=\delta_{y_{0}(t)}(\mathsf{d}y)$, where we denote by $y_{0}(t)$ the largest atom in $U_{t}$ for any $t\in I$.
@@
See for example [10, Section 4.2] for the proof of the above result
@@
It is shown in [16, Remark 1.9] that the open ASEP stationary measure depends real analytically on the jump rates $\alpha,\beta,\gamma,\delta$ and $q$ in the region of the parameter space where the stationary measure is unique. The continuity with respect to parameters $A,B,C,D$ then follows from [4, equation (2.4)]:
$\alpha=\frac{1-q}{(1+C)(1+D)},\quad\beta=\frac{1-q}{(1+A)(1+B)},\quad\gamma=\frac{-(1-q)CD}{(1+C)(1+D)},\quad\delta=\frac{-(1-q)AB}{(1+A)(1+B)}.$
The following result is known as the `stochastic sandwiching' of open ASEP stationary measure, which is first introduced by [17]:
@@
As a corollary, for any $t_{1},\dots,t_{n}\geq 1$, we have $\mathbb{E}_{\mu^{\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right]\leq\mathbb{E}_{\mu^{\prime\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right]$.
2.2. Proof of Theorem 1.2 and Theorem 1.3
@@
For reasons that will become clear later, we first prove the results modulo a particle-hole duality. The main parts of these theorems (modulo the particle-hole duality) are stated in the following result, which combines both the high and low density phases:
@@
Assume $\max(A,C)>1$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists $\varepsilon>0$ and probability measures $\eta_{m}$ on $\{0,1\}^{m}$ for $m=1,2,\dots$ such that for $1\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$, we have:
$\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right).$ (2.9)
The marginal distribution of $\eta_{m+1}$ on the first $m$ sites coincides with $\eta_{m}$.
@@
The following result provides a subtle bound on the total variation of signed measures $P_{s,t}(x,\mathsf{d}y)$:
@@
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
@@
The above result will appear again as Proposition B.1, which will be proved in Appendix B. It is worth noting that a strictly weaker version of this total variation bound has previously appeared in [10]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [10, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. The derivation of the bound here requires a delicate analysis of atom masses, see Appendix B.
@@
Step 1
In this step, under an extra assumption $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$, we prove that there exists $\varepsilon>0$ such that, for $1\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$, we have:
$\lim_{n\rightarrow\infty}\mathbb{E}_{\mu_{m,n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right).$ (2.11)
@@
We choose $\varepsilon>0$ and $I=[1,1+\varepsilon)$ according to Theorem 2.5 and Proposition 2.12. By Theorem 2.5, we have that for any $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
@@
where we write:
$\begin{split}\Gamma_{1}^{\prime}&:=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}^{\prime}&:=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.13)
@@
$\begin{split}\Gamma_{1}&:=\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ &=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}^{m}}(2+2y_{0}(1))^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}&:=\pi_{1}\left(\{y_{0}(1)\}\right)=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\{y_{0}(1)\}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.14)
@@
Hence we need to bound three quantities $|\Gamma_{1}/\Gamma_{2}|$, $|\Gamma_{1}^{\prime}-\Gamma_{1}|$ and $|\Gamma_{2}^{\prime}-\Gamma_{2}|$. We first recall that:
@@
$r:=\sup_{t\in I}\left(1+t+2\sqrt{t}y_{0}(t)\right).$
In view of (2.16) and (2.17), we have:
$\left\lvert\frac{\Gamma_{1}}{\Gamma_{2}}\right\lvert\leq r^{m}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.18)
@@
$\Gamma_{1}^{\prime}-\Gamma_{1}=\\ \frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\left(U_{1}\setminus\{y_{0}(1)\}\right)\times U_{t_{1}}\times\dots\times U_{t_{m}}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right).$ (2.19)
@@
$\left\lVert\pi_{1,t_{1},\dots,t_{m}}\right\rVert_{TV}\leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{i=1}^{m}\sup_{x\in U_{t_{i-1}}}\left\lVert P_{t_{i-1},t_{i}}\left(x,\mathsf{d}y\right)\right\rVert_{TV}\leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$
@@
$0<\frac{(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})}{\left(2+2y_{0}(1)\right)^{n}}\leq\frac{(2+2y^{*}_{0}(1))^{n-m}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}<\frac{(2+2y^{*}_{0}(1))^{n}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}=\theta^{n}r^{m},$
@@
Using a similar argument as above, we conclude
$|\Gamma_{2}^{\prime}-\Gamma_{2}|\leq\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}.$ (2.21)
@@
$|\Gamma_{2}|=\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert>\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}\geq|\Gamma_{2}^{\prime}-\Gamma_{2}|.$
Combining the estimates (2.15), (2.18), (2.20) and (2.21), we get:
@@
For fixed $m$, as $n\rightarrow\infty$, the RHS above converges to 00. In view of (2.12) and the definitions of $\Gamma_{1}$ and $\Gamma_{2}$ (so that the RHS of (2.11) equals $\Gamma_{1}/\Gamma_{2}$), we conclude the proof of the convergence (2.11).
Step 1 is complete.
Step 2
In this step, under the extra assumption $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$, we prove (2.10): there exists $H>0$ depending on $A,B,C,D$ and $q$ such that for all $1\leq m\leq n$,
@@
One can choose $H=H(A,B,C,D,q)>0$ sufficiently large so that the RHS of (2.26) can be bounded above by $\theta^{n}(Hm)^{3m}$ for any $1\leq m\leq n$ and $n\geq N$. For $n<N$ one can increase the value of $H$ if necessary, so that $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)$ can still be bounded by $\theta^{n}(Hm)^{3m}$ for any $1\leq m\leq n$. We conclude the proof of (2.23). Step 2 is complete.
Step 3
In this step we show that the total variation distance bound (2.23) continues to hold without the assumption $ABCD\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$. Note that the well-definedness of the measure $\eta_{m}$ (which we concluded from Step 1) does not have any issues, since it only involves $A,B,C$ and $q$.
@@
We return to Step 3 in the proof. We consider the case $ABCD=q^{-j}$ for some $j\in\mathbb{N}_{0}$. We will use the continuity argument. We take a sequence of $D_{k}\in(-1,0)$, $k=1,2,\dots$ that converges to $D$, satisfying $ABCD_{k}\notin\{q^{-l}:l\in\mathbb{N}_{0}\}$ for all $k$. Since $(ABCD)_{\infty}=0$, we can assume that $|(ABCD_{k})_{\infty}|\leq 1$ for all $k$. Using the continuity of the stationary measure $\mu_{n}^{(A,B,C,D)}$ from Lemma 2.8, we have:

### v2

@@ 2.1 Background
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $c=q^{j}As/\sqrt{t}$ and $d=1/\left(q^{j}A\sqrt{t}\right)<1$. Only $a$ and $c$ in $\{a,b,c,d\}$ can have norm $\geq 1$. We have $c/a=q^{j}s/t\notin\{q^{l}:l\in\mathbb{Z}\}$ since $s/t\in(q,1)$. Hence assumption (2) holds.
@@
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $c=q^{j}C/\sqrt{t}$ and $d=s/(q^{j}C\sqrt{t})<1$. Only $a$ and $c$ in $\{a,b,c,d\}$ can have norm $\geq 1$. If $A<1$ then $a=A\sqrt{t}<1$ and assumption (2) vacuously holds. If $C<1$ then $c=q^{j}C/\sqrt{t}<1$ and assumption (2) vacuously holds. If $A,C\geq 1$ then $c/a=q^{j}C/(At)\notin\{q^{l}:l\in\mathbb{Z}\}$ by our assumption, hence assumption (2) holds.
We conclude the proof. ∎
@@
Assume that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Choose the interval $I=[1,1+\varepsilon)$ from Theorem 2.5. Then for any $s\leq t$ in $I$ and any $x\in U_{s}$, the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is supported on $U_{t}$. In the high density phase, we have $P_{s,t}\left(y_{0}(s),\mathsf{d}y\right)=\delta_{y_{0}(t)}(\mathsf{d}y)$, where we denote by $y_{0}(t)$ the largest atom in $U_{t}$ for any $t\in I$.
@@
The stationary measures $\mu_{n}^{(A,B,C,D)}$ and $\mu_{n}^{(C,D,A,B)}$ for open ASEP on $\{1,\dots,n\}$ are related by a transformation of occupation variables
@@
See for example [30, Section 4.2] for the proof of the above result.
@@
It is shown in [2, Remark 1.9] that the open ASEP stationary measure depends real analytically on the jump rates $\alpha,\beta,\gamma,\delta$ and $q$ in the region of the parameter space where the stationary measure is unique. The continuity with respect to parameters $A,B,C,D$ then follows from [9, equation (2.4)]:
$\alpha=\frac{1-q}{(1+C)(1+D)},\quad\beta=\frac{1-q}{(1+A)(1+B)},\quad\\ \gamma=\frac{-(1-q)CD}{(1+C)(1+D)},\quad\delta=\frac{-(1-q)AB}{(1+A)(1+B)}.$
The following result is known as the ‘stochastic sandwiching’ of open ASEP stationary measure, which is first introduced by [12]:
@@
As a corollary, for any $t_{1},\dots,t_{n}\geq 1$, we have
$\mathbb{E}_{\mu^{\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right]\leq\mathbb{E}_{\mu^{\prime\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right].$
2.2 Proof of Theorem 1.2 and Theorem 1.4
@@
For reasons that will become clear later, we first prove the results modulo a particle-hole duality. The main parts of these theorems (modulo the particle-hole duality) are stated in Theorem 2.11 below, which combines both the high and low density phases.
Definition 2.10.
Assume $\max(A,C)>1$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. We define probability measures $\eta_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ by their joint generating functions: For some $\varepsilon>0$ and for any $t_{1}\leq\dots\leq t_{m}$ within the interval $[1,1+\varepsilon)$,
$\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\\ =\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right),$ (2.9)
where we recall from Lemma 2.6 that $y_{0}(1)$ is the largest atom in the (compact) support $U_{1}$ of the Askey–Wilson signed measure $\pi_{1}(\mathsf{d}y)=\nu(\mathsf{d}y;A,B,C,D)$, and $P_{s,t}(x,\mathsf{d}y)$ was defined by (2.6). We will prove in Theorem 2.11 that there exists $\varepsilon>0$ depending on $A,B,C,D$ and $q$ such that for all $m\in\mathbb{Z}_{+}$, the probability measure $\eta_{m}$ on $\{0,1\}^{m}$ are well-defined.
@@
Assume $\max(A,C)>1$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then the probability measures $\eta_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ in Definition 2.10 are well-defined, and the marginal distribution of $\eta_{m+1}$ on the leftmost sublattice $\{1,\dots,m\}$ coincides with $\eta_{m}$.
@@
The following result provides a delicate bound of the total variation of Askey–Wilson signed measures $P_{s,t}(x,\mathsf{d}y)$:
@@
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
@@
The above result will appear again as Proposition B.1, which will be proved in Appendix B. It is worth noting that a strictly weaker version of this total variation bound has previously appeared in [30]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [30, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. The derivation of the bound here requires a delicate analysis of atom masses, see Appendix B.
@@
Step 1. In this step, under an extra assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$, we prove that there exists $\varepsilon>0$ such that, for $1\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$, we have:
$\lim_{n\rightarrow\infty}\mathbb{E}_{\mu_{m,n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\\ =\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right).$ (2.11)
@@
We choose $\varepsilon>0$ and $I=[1,1+\varepsilon)$ according to Theorem 2.5 and Proposition 2.13. By Theorem 2.5, we have that for any $1\leq m<n$ and $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
@@
where we define:
$\begin{split}\Gamma_{1}^{\prime}:=&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}^{\prime}:=&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.13)
@@
$\begin{split}\Gamma_{1}:=&\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ =&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\\ &\times\int_{\mathbb{R}^{m}}(2+2y_{0}(1))^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}:=&\pi_{1}\left(\{y_{0}(1)\}\right)=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\{y_{0}(1)\}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.14)
@@
Hence we need to bound three quantities $|\Gamma_{1}/\Gamma_{2}|$, $|\Gamma_{1}^{\prime}-\Gamma_{1}|$ and $|\Gamma_{2}^{\prime}-\Gamma_{2}|$. We first recall that $\Gamma_{1}/\Gamma_{2}$ is the RHS of (2.11), i.e.,
@@
$r:=\sup_{t\in I}\left(1+t+2\sqrt{t}y_{0}(t)\right)$
In view of (2.16) and (2.17), and noting that $12+2y_{0}(t)\geq 4>$ 1, we have:
$\left\lvert\frac{\Gamma_{1}}{\Gamma_{2}}\right\lvert\leq\left(\frac{r}{2+2y_{0}(1)}\right)^{m}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}\leq r^{m}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.18)
@@
$\Gamma_{1}^{\prime}-\Gamma_{1}=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\left(U_{1}\setminus\{y_{0}(1)\}\right)\times U_{t_{1}}\times\dots\times U_{t_{m}}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\\ \times\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right).$ (2.19)
@@
$\left\lVert\pi_{1,t_{1},\dots,t_{m}}\right\rVert_{TV}\leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{i=1}^{m}\sup_{x\in U_{t_{i-1}}}\left\lVert P_{t_{i-1},t_{i}}\left(x,\mathsf{d}y\right)\right\rVert_{TV}\\ \leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$
@@
$0<\frac{(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})}{\left(2+2y_{0}(1)\right)^{n}}\\ \leq\frac{(2+2y^{*}_{0}(1))^{n-m}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}<\frac{(2+2y^{*}_{0}(1))^{n}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}=\theta^{n}r^{m},$
@@
Using a similar but simpler argument to the one above, we conclude
$|\Gamma_{2}^{\prime}-\Gamma_{2}|\leq\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}.$
@@
$|\Gamma_{2}|=\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert>\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}\geq|\Gamma_{2}^{\prime}-\Gamma_{2}|.$ (2.21)
Combining the estimates (2.15), (2.18), (2.20) and (2.21), we get that for any $n\geq N$, $1\leq m<n$ and $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
@@
For fixed $m$, as $n\rightarrow\infty$, the RHS above converges to 00. In view of (2.12) and the definitions of $\Gamma_{1}$ and $\Gamma_{2}$ (so that the RHS of (2.11) equals $\Gamma_{1}/\Gamma_{2}$), we conclude the proof of the convergence (2.11). Step 1 is complete.
Step 2. In this step, under the extra assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$, we prove (2.10): there exists $H>0$ depending on $A,B,C,D$ and $q$ such that for all $1\leq m\leq n$,
@@
One can choose $H=H(A,B,C,D,q)>0$ sufficiently large so that the RHS of (2.26) can be bounded above by $\theta^{n}(Hm)^{3m}$ for any $n\geq N$ and $1\leq m<n$. For the cases when $n<N$ or $m=n$, one can increase the value of $H$ if necessary, so that $\theta^{n}(Hm)^{3m}\geq 1$, which is greater or equal to $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq 1$. We conclude the proof of (2.23). Step 2 is complete.
Step 3. In this step we show that the total variation distance bound (2.23) continues to hold without the assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$. Note that the well-definedness of the measure $\eta_{m}$ (which we concluded from Step 1) does not have any issues, since it only involves $A,B,C$ and $q$.
@@
We return to Step 3 in the proof. We consider the case $q^{j}ABCD=1$ for some $j\in\mathbb{N}_{0}$. We will use the continuity argument. We take a sequence of $D_{k}\in(-1,0)$, $k=1,2,\dots$ that converges to $D$, satisfying $q^{l}ABCD_{k}\neq 1$ for all $l\in\mathbb{N}_{0}$ and for all $k$. Since $(ABCD)_{\infty}=0$, we can assume that $|(ABCD_{k})_{\infty}|\leq 1$ for all $k$. Using the continuity of the stationary measure $\mu_{n}^{(A,B,C,D)}$ from Lemma 2.8, we have: